OpenStudy (anonymous):

What is the ratio of a to b? 8/15=4a/a+b

OpenStudy (tkhunny):

Do you mean \(\dfrac{8}{15} = \dfrac{4a}{a+b}\)? That is NOT what you have written, but it may be what you intended. Please confirm.

OpenStudy (anonymous):

8/15=4a/a+b 8(a+b) = 4a (15) 8a+8b = 60a transpose

OpenStudy (anonymous):

yes that what i confirm

OpenStudy (anonymous):

You can solve for a different ways. I think the easiest is to try to find a by using the numerator first, so 4a = 8, divide both sides by 4, a = 2, now that we now a = 2 we can solve the denominator for b 2 + b = 15, subtract both sides by 2, b = 13 Now that you know a = 2 and b = 13 you can make the ratio of a to b 2:13